OpenStudy (anonymous):

write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0. -√5,5√5

OpenStudy (anonymous):

\[(x+\sqrt{5})(x-5\sqrt{5})\] \[x ^{2}+4\sqrt{5}-25=0\]